OpenStudy (keana):

2x^2 + 7x + 3 = 0. isn't it -1/2, -3

OpenStudy (anonymous):

2x^2+6x+x+3=0 2x(x+3)+1(x+3)=0 (2x+1)(x+3)=0 x=-1/2,-3

OpenStudy (anonymous):

got it?